# Supersingular curves via the Shimura–Taniyama method

### v1

@@ 4. Cyclic Examples Genus 6 through 10
Our program computed the following data for curves $X$, of genus $6\leq g\leq 10$, that admit a degree $m$ cyclic cover $\phi:X\to\mathbf{P}^{1}$ branched at three points: the ramification type $z$; the possible inertia types $a$, up to equivalence; and the odd primes $p$ with $\mathrm{gcd}(p,m)=1$ such that $X$ is supersingular. The results are listed in Tables 1-5
@@ 5. Non-Cyclic Examples
Our program identified $17$ possible ramification types $z=[c_{0},c_{1},c_{\infty},s]$ for curves $X$ admitting an abelian cover $\phi:X\to\mathbf{P}^{1}$ branched at three points with $s>1$ such that the genus $g$ satisfies $6\leq g\leq 10$.
@@
In this section, we analyze the 14 valid cases and prove a general result about a class of examples which occur for every genus.
Throughout, we use the following notation for quotients. Suppose $g\in G$ has prime order $\ell$. Let $X_{g}=X/\langle g\rangle$ be the quotient curve and $\phi_{g}:X\to X_{g}$ be the quotient cover. For $b\in\{0,1,\infty\}$, the inertia group $I_{b}$ is independent of the ramification point $\eta\in X$ above $b$ since $G$ is abelian. If $g\in I_{b}$, then $\phi_{g}$ is ramified with index $\ell$ at $\eta$. If $g\not\in I_{b}$, then $\phi_{g}$ is not ramified at $\eta$.
@@ 5.1. A useful lemma
Let $r\geq 2$. If $z=[2r,2r,2,2]$, then $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$, where $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ is an abelian cover branched at $B=\{0,1,\infty\}$ for $i=1,2$. If $r$ is odd and $i=1,2$, or if $r$ is even and $i=1$, then $\phi_{i}$ has degree $2r$ with inertia type $a=(1,r,r-1)$. If $r$ is even, then $\phi_{0}$ has Galois group $\mathbf{Z}/r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ with inertia type $a=((1,0),(-1,-1),(0,1))$.
@@
Let $K=\langle(r,0),(0,1)\rangle$. Note that $X/K$ also has genus 00, since it is a quotient of $X_{a_{\infty}}$. Write $X_{0}=X/\langle(r,0)\rangle$ and $X_{1}=X/\langle(r,1)\rangle$. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$.
@@
Suppose $r$ is odd. By the Riemann–Hurwitz formula, $X_{0}$ and $X_{1}$ both have genus $(r-1)/2$. The quotient group $\overline{G}_{0}=G/\langle(r,0)\rangle$ is cyclic of order $2r$ with generator $\tau_{0}=(1,1)$. Then $a_{1}=-\tau_{0}$, $a_{\infty}\equiv r\tau_{0}\bmod(r,0)$, and $a_{0}=(r+1)\tau_{0}$. Thus the reduction of the inertia type in $\overline{G}_{0}$ is $(r+1,r,-1)\approx(1,r,r-1)$. The quotient group $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,0)=a_{0}$. We see that $a_{\infty}\equiv r\tau_{1}$ and $a_{1}\equiv(r-1)\tau_{1}$. Thus the $\mathbf{Z}/2r\mathbf{Z}$-cover $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ has inertia type $(1,r,r-1)$ for $i=1,2$.
Suppose $r$ is even. Then $\phi_{(r,1)}$ is unramified and $X_{1}$ has genus $r/2$; also $\phi_{(r,0)}$ is ramified at four points and $X_{0}$ has genus $(r/2)-1$. Also $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,1)=-a_{1}$. By a similar argument, $\phi_{1}$ has inertia type $(1,r,r-1)$. The inertia type for $\phi_{0}$ is found by working in $G/\langle(r,0)\rangle$. ∎
@@
Recall that $\mathbf{S}_{g}$ is the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. Using Lemma 5.2, for $r\geq 2$, one can show that the ramification type $z=[2r,2r,2,2]$ yields a curve $X$ of genus $r-1$ which is supersingular for all primes $p\equiv-1\bmod 2r$. This shows that the natural density of $\mathbf{S}_{g}$ is positive. We omit the details because this can also be proven using earlier work from [Yui80]. An unpublished study of supersingular diagonal curves (analyzed as quotients of the Fermat curve) also shows the density is positive [CMS].
@@ 5.2. Non-Cyclic Examples: Genus 6 through 10
The tables below contain the $14$ non-cyclic cases for $6\leq g\leq 10$. For most ramification types $z$, there is only one equivalence class of inertia type. In each case, we computed the decomposition of the Jacobian. The second column, titled decomposition, gives the data $(m_{i},\alpha_{i})$ indexed by $i\in I$. This indicates that there is an isogeny $\mathbf{Jac}(X)\sim\prod_{i\in I}\mathbf{Jac}(Z_{i})$, where there is a cyclic degree $m_{i}$ cover $\phi_{i}:Z_{i}\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ with inertia type $\alpha_{i}$. We use an addition sign to denote a product of Jacobians, an exponent to denote a product with multiplicity, and a subtraction sign to denote a quotient. From the decomposition, we used the Shimura–Taniyama method for cyclic covers implemented in our program to determine the congruence conditions on $p$, with $\mathrm{gcd}(p,|G|)=1$, for which $X$ is supersingular.
@@
There are two ramification types for genus six curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
@@
There are three ramification types for genus seven curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
@@
There are two ramification types for genus eight curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
@@
There are three ramification types for genus nine curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
@@
There are four ramification types for genus ten curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
@@
By [Igu60, Section 8, page 645], a curve of genus $2$ with an automorphism of order $6$ is isomorphic to $y^{2}=x(x-1)(x+1)(x-2)(x-(1/2))$. By [IKO86, Proposition 1.11], this curve is supersingular if and only if $p\equiv 5\bmod 6$. Applying Theorem 2.12 twice (to the quotient of $X_{2,4}$ by the Klein four action and to the quotient of $X$ by $H$) shows that $X$ is supersingular if and only if $p\equiv 5\bmod 6$.
If $z=[9,9,9,3]$, then $G=\mathbf{Z}/9\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(1,1),(7,2))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotient $X_{(3,0)}$ and the quotient $X/H$ are both elliptic curves, having an automorphism of order $3$. Each of $X_{(3,1)}$, $X_{(3,2)}$ and $X_{(0,1)}$ has genus $4$, and is a $\mathbf{Z}/9\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with inertia type $\bar{a}\approx(1,1,7)$. Then use Theorem 2.12.
If $z=[12,12,3,3]$, then $G=\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(11,2),(0,1))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotients $X_{(0,1)}$ and $X_{H}$ both have genus 00. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{(4,0)})\times\mathbf{Jac}(X_{(4,1)})\times\mathbf{Jac}(X_{(4,2)})$. Then $X_{(4,0)}$ is a $\mathbf{Z}/4\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with $a=((1,0),(3,2),(0,1))$, which is equivalent to a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(1,3,8)$. For $X_{(4,1)}$, we note that $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,1)\rangle$ is generated by $\tau=(1,2)$. So $X_{(4,1)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,3\tau,4\tau)\approx(1,3,8)$. Similarly, $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,2)\rangle$ is generated by $\tau=(1,1)$ and $X_{(4,2)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,11\tau,8\tau)\approx(1,4,7)$.
@@ 6. Conclusions
For $5\leq g\leq 10$: the primes $p$ for which there exists a supersingular curve $X$ of genus $g$ having an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ are:
@@
    for $g=5$, any $p$ such that $p\equiv QNR\bmod 11$ and/or
    $p\equiv 7,11,19,23,31,41,47,59,71,79,83,89,91,101,103,107,119\bmod 120.$
    Thus $\delta_{5}=25/32$.
@@
    for $g=10$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv QNR\bmod 11$; and $p\not\equiv 1\bmod 40$. Thus $\delta_{10}=63/64$.

### v2

@@ 4. Cyclic Examples Genus 6 through 10
Our program computed this data for curves $X$ over $k$, of genus $6\leq g\leq 10$, that occur as a degree $m$ cyclic cover $\phi:X\to\mathbf{P}^{1}$ branched at three points: the ramification types $z$; the inertia types $a$, up to equivalence; and the odd primes $p$ with $\mathrm{gcd}(p,m)=1$ for which the reduction of $X$ modulo $p$ is supersingular. The results are in Tables 1-5.
@@ 5. Non-Cyclic Examples
Our program identified $14$ ramification types $z=[c_{0},c_{1},c_{\infty},s]$, with $s>1$, for which there is a valid inertia type, for an abelian cover $\phi:X\to\mathbf{P}^{1}$ branched at three points such that the genus $g$ of $X$ satisfies $6\leq g\leq 10$. In this section, we analyze the supersingular primes $p$ for these covers, assuming throughout that $p\nmid c_{0}s$. We also prove a general result about a class of examples which occur for every genus.
@@
Here is some notation. If $g\in G$ has prime order $\ell$, let $X_{g}=X/\langle g\rangle$ be the quotient curve and $\phi_{g}:X\to X_{g}$ be the quotient cover. For $b\in\{0,1,\infty\}$, the inertia group $I_{b}$ is independent of the ramification point $\eta\in X$ above $b$ since $G$ is abelian. If $g\in I_{b}$, then $\phi_{g}$ is ramified with index $\ell$ at $\eta$. If $g\not\in I_{b}$, then $\phi_{g}$ is not ramified at $\eta$.
@@ 5.1. A useful lemma
Let $r\geq 2$. If $z=[2r,2r,2,2]$, then $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{1})\times\mathbf{Jac}(X_{2})$, where $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ is an abelian cover branched at $B=\{0,1,\infty\}$ for $i=1,2$. If $r$ is odd and $i=1,2$, or if $r$ is even and $i=1$, then $\phi_{i}$ has group $\mathbf{Z}/2r\mathbf{Z}$ with $a=(1,r,r-1)$. If $r$ is even, then $\phi_{2}$ has group $\mathbf{Z}/r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ with $a=((1,0),(-1,-1),(0,1))$.
@@
Let $K=\langle(r,0),(0,1)\rangle$. Since $X/K$ is a quotient of $X_{a_{\infty}}$, it also has genus 00. Write $X_{1}=X/\langle(r,1)\rangle$ and $X_{2}=X/\langle(r,0)\rangle$. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{1})\times\mathbf{Jac}(X_{2})$.
@@
Suppose $r$ is odd. Then $X_{1}$ and $X_{2}$ both have genus $(r-1)/2$. The quotient group $\overline{G}_{2}=G/\langle(r,0)\rangle$ is cyclic of order $2r$ with generator $\tau_{0}=(1,1)$. Then $a_{1}=-\tau_{0}$, $a_{\infty}\equiv r\tau_{0}\bmod(r,0)$, and $a_{0}=(r+1)\tau_{0}$. Thus the reduction of the inertia type in $\overline{G}_{2}$ is $(r+1,r,-1)\approx(1,r,r-1)$. The quotient group $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,0)=a_{0}$. We see that $a_{\infty}\equiv r\tau_{1}$ and $a_{1}\equiv(r-1)\tau_{1}$. Thus the $\mathbf{Z}/2r\mathbf{Z}$-cover $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ has inertia type $(1,r,r-1)$ for $i=1,2$.
Suppose $r$ is even. Then $\phi_{(r,1)}$ is unramified and $X_{1}$ has genus $r/2$; also $\phi_{(r,0)}$ is ramified at four points and $X_{2}$ has genus $(r/2)-1$. Also $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,1)=-a_{1}$. By a similar argument, $\phi_{1}$ has inertia type $(1,r,r-1)$. The inertia type for $\phi_{2}$ is found by working in $G/\langle(r,0)\rangle$. ∎
@@
By Lemma 5.2, for $r\geq 2$, the ramification type $z=[2r,2r,2,2]$ yields a curve $X$ of genus $r-1$ which is supersingular for all primes $p\equiv-1\bmod 2r$. This shows that the natural density of $\mathbf{S}_{g}$ is positive. We omit the details because this can also be proven using [Yui80] or (unpublished) work in [CMS].
@@ 5.2. Non-Cyclic Examples: Genus 6 through 10
The tables below contain the following information for the $14$ non-cyclic cases for $6\leq g\leq 10$: the ramification type $z$; the decomposition of the Jacobian; the congruence conditions on $p$, assuming $\mathrm{gcd}(p,|G|)=1$, for which the curve $X$ is supersingular.
The decomposition of the Jacobian is given by the data $(m_{i},\alpha_{i})$ indexed by $i\in I$, meaning that there is an isogeny $\mathbf{Jac}(X)\sim\prod_{i\in I}\mathbf{Jac}(Z_{i})$, where there is a cyclic degree $m_{i}$ cover $\phi_{i}:Z_{i}\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ with inertia type $\alpha_{i}$. We use an addition sign to denote a product of Jacobians, an exponent to denote a product with multiplicity, and a subtraction sign to denote a quotient.
@@
Here are the supersingular primes for genus six curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
@@
Here are the supersingular primes for genus seven curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
@@
Here are the supersingular primes for genus eight curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
@@
Here are the supersingular primes for genus nine curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
@@
Here are the supersingular primes for genus ten curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
@@
By [Igu60, Section 8, page 645], a curve of genus $2$ with an automorphism of order $6$ is isomorphic to $y^{2}=x(x-1)(x+1)(x-2)(x-(1/2))$. By [IKO86, Proposition 1.11], this curve is supersingular if and only if $p\equiv 5\bmod 6$. Applying Theorem 2.12 twice shows that $X$ is supersingular if and only if $p\equiv 5\bmod 6$.
If $z=[9,9,9,3]$, then $G=\mathbf{Z}/9\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(1,1),(7,2))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotients $X_{(3,0)}$ and $X/H$ are both elliptic curves, having an automorphism of order $3$. Each of $X_{(3,1)}$, $X_{(3,2)}$ and $X_{(0,1)}$ has genus $4$, and is a $\mathbf{Z}/9\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with inertia type $\bar{a}\approx(1,1,7)$. Then use Theorem 2.12.
If $z=[12,12,3,3]$, then $G=\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(11,2),(0,1))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotients $X_{(0,1)}$ and $X_{H}$ both have genus 00. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{(4,0)})\times\mathbf{Jac}(X_{(4,1)})\times\mathbf{Jac}(X_{(4,2)})$. Then $X_{(4,0)}$ is a $\mathbf{Z}/4\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with $a=((1,0),(3,2),(0,1))$, which is equivalent to a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(1,3,8)$. Because $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,1)\rangle$ is generated by $\tau=(1,2)$, $X_{(4,1)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,3\tau,4\tau)\approx(1,3,8)$. Similarly, $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,2)\rangle$ is generated by $\tau=(1,1)$ and $X_{(4,2)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,11\tau,8\tau)\approx(1,4,7)$.
@@ 6. Conclusions
The primes $p$ for which there exists a supersingular curve $X$ of genus $5\leq g\leq 10$ having an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ are:
@@
    for $g=5$, any $p\equiv 7,11,19,23,31,41,47,59,71,79,83,89,91,101,103,107,119\bmod 120$ and/or $p\equiv QNR\bmod 11$. Thus $\delta_{5}=25/32$.
@@
    for $g=10$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv QNR\bmod 11$; $p\not\equiv 1,11,31\bmod 40$; and $p\equiv 3,19,27\pmod{28}$. Thus $\delta_{10}=31/32$.